rolling a die and drawing a card from a deck find P (even and black)
flip a coin and draw from a deck of cards find p (heads and 7)

Answer by stanbon(75887)

rolling a die and drawing a card from a deck find P (even and black)
The events are independent, so
P(even and black) = P(even)*P(black) = (1/2)(1/2) = 1/4
---------------------------------------------------------------

flip a coin and draw from a deck of cards find p (heads and 7)
The events are independent, so
P(heads and 7) = P(heads)*P(7) = (1/2)*(4/52) = 1/26
==============================================================
Cheers,
Stan H.